OpenStudy (anonymous):

4(2x -3) + 8(x - 4) = 2(2x + 6)

OpenStudy (anonymous):

ok first you distribute. do you know how to do that?

OpenStudy (anonymous):

ive already tried and i got x=12 but its a fraction anser

OpenStudy (anonymous):

its 16x-12+8x-4= 4x+12 now you add the same variables

OpenStudy (anonymous):

so its 24x-16=4x+12 now you subtract for and add 16

OpenStudy (anonymous):

its not 12

OpenStudy (anonymous):

it hasa to be a fraction

OpenStudy (anonymous):

or 4

OpenStudy (anonymous):

the answer is x= (7/5)